Question No. 4

 

The Gauteng chamber of business conducted a survey amongst 17 furniture retailers to identify the percentage of bad debts in each company’s debtors’ book. The bad debts percentages are as follows:

 

2.2, 4.7, 6.3, 5.8, 5.7, 7.2, 2.6, 2.4, 6.1, 6.8, 2.2, 5.7, 3.4, 6.6, 1.8, 4.4, 5.4

 

Calculate the Pearson Coefficient of skewness coefficient for percentage of bad debts. Is the data skewed?


1
Expert's answer
2021-05-28T09:28:54-0400

"Sk=\frac{3(\bar{X}-Md)}{s}"

Sk is coefficient of skewness

Md is median

s is standard deviation

1.8

2.2

2.2

2.4

2.6

3.4

4.4

4.7

5.4

5.7

5.7

5.8

6.1

6.3

6.6

6.8

7.2 Md 5.4 s 1.851669 "\bar{X}=4.664705882"

"Sk=\frac{3(4.66471-5.4)}{1.851669}=-1.1913"

The data is negatively skewed due to the negative sign. The rule of thumb states that the coefficient of skewness should be between -1 and +1. thus, -1.1913 shows high negatively skewed data.